# Differentially Private Synthetic Data via Foundation Model APIs 1: Images

### v2

@@ Abstract
${}^{\dagger}$ Main updates in arXiv V2: (1) Fixing typos in the convergence analysis (Apps. D and E). (2) Adding additional experiments in Apps. M, L, J, I and 5.1.2. (3) Polishing the writing.
@@ 1 Introduction
Figure 3: Generated samples on CIFAR10 with $\left(0.67,10^{-5}\right)$-DP. Each row corresponds to one class. FID=7.87. See App. I for real and generated images side-by-side.
Figure 3: Generated samples on CIFAR10 with $\left(0.67,10^{-5}\right)$-DP. Each row corresponds to one class. FID=7.87. See App. I for real and generated images side-by-side.
@@
(2) New framework (§ 4). We propose Private Evolution (PE) algorithm for achieving our goal (Fig. 2). We consider using 2 popular APIs: random generation and sample variation (i.e., generating a sample similar to the given one).Footnotes 4 and 5 The key idea is to iteratively use private samples to vote for the most similar samples generated from the blackbox model and ask the blackbox models to generate more of those similar samples. We theoretically prove that the distribution of the generated samples from PE will converge to the private distribution under some modeling assumptions (App. D). PE only requires (existing) APIs of the foundation models, and does not need any model training.
@@ 3.2 Problem Formulation
DP Wasserstein Approximation (DPWA). Given a private dataset $S_{\mathrm{priv}}=\left\{x_{i}:i\in[N_{\mathrm{priv}}]\right\}$ with $N_{\mathrm{priv}}$ samples (e.g., images), a distance function $d(\cdot,\cdot)$ between samples and some $p\geq 1$, the goal is to design an $(\epsilon,\delta)$-DP algorithm $\mathcal{M}$ that outputs a synthetic dataset $S_{\mathrm{syn}}=\left\{x_{i}^{\prime}:i\in[N_{\mathrm{syn}}]\right\}$ with $N_{\mathrm{syn}}$ samples (as a multiset) whose distance to $S_{\mathrm{priv}}$, $W_{p}(S_{\mathrm{priv}},S_{\mathrm{syn}})$, is minimized. Here $W_{p}$ is the Wasserstein $p$-distance w.r.t. the distance function $d(\cdot,\cdot)$ (see App. A for the definition).
@@
When the distance function $d(\cdot,\cdot)$ is $\ell_{2}$ (in the sample space), DPWA is closely related to DP Clustering (Ghazi et al., 2020; Su et al., 2016; Balcan et al., 2017) and DP Heatmaps (Ghazi et al., 2022). But a direct application of them does not work in our setting; see App. G for more discussions.
@@ 4 Private Evolution (PE)
Inspired by evolutionary algorithms (EA) (Davis, 1987) (App. B), we propose Private Evolution (PE) framework for generating DP synthetic data via APIs. See Fig. 2 for the intuition behind PE. The complete algorithm is in Alg. 1. Below, we discuss the components in detail.
@@
1 $S_{1}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}\right)$ for $t\leftarrow 1,\ldots,T$ do
       $histogram_{t}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(S_{\mathrm{priv}},S_{t},\sigma,H\right)$   // See Alg. 2
       $\mathcal{P}_{t}\leftarrow histogram_{t}/\mathrm{sum}(histogram_{t})$   // $\mathcal{P}_{t}$ is a distribution on $S_{t}$
       $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}$ samples with replacement from $\mathcal{P}_{t}$   // $S_{t}^{\prime}$ is a multiset
@@ 2       $S_{t+1}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
1 $histogram\leftarrow[0,\ldots,0]$ for $x_{\mathrm{priv}}\in S_{\mathrm{priv}}$ do
2       $i=\arg\min_{j\in\left[n\right]}d{}\left(x_{\mathrm{priv}},z_{j}\right)$ $histogram[i]\leftarrow histogram[i]+1$
$histogram\leftarrow histogram+\mathcal{N}\left(0,\sigma I_{n}\right)$   // Add noise to ensure DP
$histogram\leftarrow\max\left(histogram-H,0\right)$   // ‘max’, ‘-’ are element-wise
@@ 1 $histogram\leftarrow[0,\ldots,0]$ for $x_{\mathrm{priv}}\in S_{\mathrm{priv}}$ do
In summary, we called the above fitness function DP Nearest Neighbors Histogram. Note that it is not the traditional “histogram” on a continuous space that requires binning. Instead, it is a histogram built on the generated samples: the value of $i$-th bin means the (privatized) number of private samples whose nearest neighbor among the generated samples is the $i$-th sample.
@@
Please see App. D for the convergence analysis of PE.
@@ 4.3 Privacy Analysis
Unlike the analysis DP-SGD which requires complicated DP composition theorems (e.g., Gopi et al. (2021); Mironov (2017)) due to subsampling, PE does not have subsampling steps and therefore the privacy analysis is rather straightforward. The DP guarantee of the unconditional version of PE (Alg. 1) can be reasoned as follows:
@@
For the conditional version of PE (Alg. 3), since it does the unconditional version of PE for each class separately (discussed in § 4.1), adding or removing one sample will only influence the generation results of one class. For that class, the impact due to the added/removed sample is also bounded, as seen in the privacy analysis above. Therefore, Alg. 3 is also DP. In fact, we can show that the privacy guarantee of Alg. 3 is the same as Alg. 1. Please refer to App. C for more details.
@@ 5 Experiments
In § 5.1, we compare PE with SOTA training-based methods on standard benchmarks to understand its promise and limitations. In § 5.2, we present proof-of-concept experiments to show how PE can utilize the power of large foundation models. We did (limited) hyper-parameter tunings in the above experiments; following prior DP synthetic data work (Yu et al., 2021; Ghalebikesabi et al., 2023), we ignore the privacy cost of hyper-parameter tuning. However, as we will see in the ablation studies (§ 5.3 and M), PE stably outperforms SOTA across a wide range of hyper-parameters, and the results can be further improved with better hyper-parameters than what we used. Detailed hyper-parameter settings and more results such as generated samples and their nearest images in the private dataset are in Apps. I, J and K.
@@ 5.1 Comparisons to State-of-the-Art
Public information. We use standard benchmarks (Ghalebikesabi et al., 2023) which treat ImageNet (Deng et al., 2009) as public data. For fair comparisons, we only use ImageNet as public information in PE: (1) Pre-trained model. Unlike the SOTA (Ghalebikesabi et al., 2023) which trains customized diffusion models, we simply use public ImageNet pre-trained diffusion models (pure image models without text prompts) (Nichol & Dhariwal, 2021). RANDOM_API and VARIATION_API are implemented using the same pre-trained model (see App. I). (2) Embedding (Eq. 1). We use ImageNet inception embedding (Szegedy et al., 2016). PE is not sensitive to embedding choice though and we get good results even with CLIP embeddings (Fig. 41 in App. M).
@@
Outline. We test PE on private datasets that are either similar to or differ a lot from ImageNet in § 5.1.1 and 5.1.2. We demonstrate that PE can generate an unlimited number of useful samples in § 5.1.3. We show that PE is computationally cheaper than train-based methods in App. O.
@@ 5.1.2 Large Distribution Shift (ImageNet $\rightarrow$ Camelyon17)
Figure 6: Real and generated images from Camelyon17. More in App. J.
@@
Why PE works under large distribution shifts. Even though the diffusion model is trained on natural images, the support of the generated distribution spans the entire sample space. PE is effective in guiding the model to generate samples from the region that is low-density in the original pre-trained distribution but high-density in the private distribution. See App. J for more explanations.
Limitation. These results demonstrate the effectiveness of PE. But when public models that are similar to private data are not available and when there is enough private data, the traditional training-based methods are still more promising at this point if the privacy-utility trade-off is the only goal. However, given the benefit of API-only assumption and the non-trivial results that PE already got, it is worth further exploiting the potential of PE in future work. Indeed, we expect these results can be improved with further refinement of PE (App. M).
@@ 5.2 More Challenging Benchmarks with Large Foundation Models
Data. Ideally we want to experiment with a dataset that has no overlap with Stable Diffusion’s training data. [^7] The training set of Stable Diffusion is public. However, it is hard to check if a public image or its variants (e.g., cropped, scaled) have been used to produce images in it. Therefore, we resort to our own private data. We take the safest approach: we construct two datasets with photos of the author’s two cats that have never been posted online. Each dataset has 100 512x512 images. Such high-resolution datasets with a small number of samples represent a common need in practice (e.g., in health care), but are challenging for DP synthetic data: to the best of our knowledge, no prior training-based methods have reported results on datasets with a similar resolution or number of samples. The dataset is released at https://github.com/microsoft/DPSDA as a new benchmark. See App. K for all images.
API implementation. We use off-the-shelf Stable Diffusion APIs (see App. K).
Results. We run PE for these two datasets with the same hyperparameters. Fig. 8 show examples of generated images for each of the cat datasets. We can see that Private Evolution correctly captures the key characteristics of these two cats. See App. K for all generated images.
@@
Figure 8: Real & generated images from Cat Cookie (left) and Cat Doudou (right). More in App. K.
@@ 5.3 Ablation Studies
Pre-trained network. Fig. 7 shows the results with two different ImageNet pre-trained networks: one is larger (270M) with ImageNet class labels as input; the other one is smaller (100M) without label input (see App. H for implementation details). In all experiments in § 5.1, we used the 270M network. Two takeaways are: (1) The 270M network (trained on the same dataset) improves the results. This is expected as larger and more powerful models can learn public distributions better. This suggests the potential of PE with future foundation models with growing capabilities. (2) Even with a relatively weak model (100M), PE can still obtain good results that beat the baselines (though with a slower convergence speed), suggesting the effectiveness of PE.
More ablation studies on the lookahead degree $k$, the number of generated samples $N_{\mathrm{syn}}$, the threshold $H$, and the embedding network are in App. M, where we see that PE obtains good results across a wide range of hyper-parameters, and the main results can be improved with better hyper-parameters.
@@ Acknowledgement
${}^{\dagger}$ This paper is the full version of our previous workshop paper (Lin et al., 2023a).
Appendix A Definition of Wasserstein Distance
@@
Appendix B A Brief Introduction to Evolutionary Algorithms
@@
Appendix C More Details on Private Evolution
@@
1 $S_{\mathrm{syn}}\leftarrow\emptyset$ for $c\in C$ do
2       $private\_samples\leftarrow\left\{x_{i}|(x_{i},y_{i})\in S_{\mathrm{priv}}\text{ and }y_{i}=c\right\}$ $S_{1}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}/\left|C\right|\right)$ for $t\leftarrow 1,\ldots,T$ do
             $histogram_{t}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(private\_samples,S_{t},\sigma,H\right)$   // See Alg. 2
             $\mathcal{P}_{t}\leftarrow histogram_{t}/\mathrm{sum}(histogram_{t})$   // $\mathcal{P}_{t}$ is a distribution on $S_{t}$
             $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}/|C|$ samples with replacement from $\mathcal{P}_{t}$   // $S_{t}^{\prime}$ is a multiset
@@ 3             $S_{t+1}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
Privacy analysis. For ease of understanding the privacy guarantee of Alg. 3, we can consider a modified version of Alg. 3 as in Alg. 4, where we switch the order of the two for loops over $t$ and $c$. Apparently, this modified algorithm gives the same outcome as Alg. 3. The only lines that touch private data are Algs. 4 and 4. The input to these lines is the entire private dataset, and the output is a histogram with size $N_{\mathrm{syn}}$. Same as step 1 in the analysis of Alg. 1 (§ 4.3), each private sample only contributes one vote in the histogram. If we add or remove one sample, the resulting histogram will change by 1 in the $\ell_{2}$ norm. Therefore, the sensitivity of these lines is 1. The following privacy analysis follows exactly the same as steps 2-5 in § 4.3, and therefore, the privacy guarantee of Alg. 3 is the same as Alg. 1.
@@
1 $S_{\mathrm{syn}}\leftarrow\emptyset$ $S_{1}^{c}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}/\left|C\right|\right)$ for each $c\in C$ $private\_samples^{c}\leftarrow\left\{x_{i}|(x_{i},y_{i})\in S_{\mathrm{priv}}\text{ and }y_{i}=c\right\}$ for each $c\in C$ for $t\leftarrow 1,\ldots,T$ do

### v3

@@ Abstract
Main updates in arXiv V3: (1) Merging updates from the ICLR 2024 camera ready version, including adding more related work in App. A. (2) Fixing typos in some hyper-parameters in Apps. J and K.
@@ 1 Introduction
Figure 3: Generated samples on CIFAR10 with $\left(0.67,10^{-5}\right)$-DP. Each row corresponds to one class. FID=7.87. See App. J for real and generated images side-by-side.
@@
(2) New framework (§ 4). We propose Private Evolution (PE) algorithm for achieving our goal (Fig. 2). We consider using 2 popular APIs: random generation and sample variation (i.e., generating a sample similar to the given one).Footnotes 4 and 5 The key idea is to iteratively use private samples to vote for the most similar samples generated from the blackbox model and ask the blackbox models to generate more of those similar samples. We theoretically prove that the distribution of the generated samples from PE will converge to the private distribution under some modeling assumptions (App. E). PE only requires (existing) APIs of the foundation models, and does not need any model training.
@@ 3.2 Problem Formulation
DP Wasserstein Approximation (DPWA). Given a private dataset $S_{\mathrm{priv}}=\left\{x_{i}:i\in[N_{\mathrm{priv}}]\right\}$ with $N_{\mathrm{priv}}$ samples (e.g., images), a distance function $d(\cdot,\cdot)$ between samples and some $p\geq 1$, the goal is to design an $(\epsilon,\delta)$-DP algorithm $\mathcal{M}$ that outputs a synthetic dataset $S_{\mathrm{syn}}=\left\{x_{i}^{\prime}:i\in[N_{\mathrm{syn}}]\right\}$ with $N_{\mathrm{syn}}$ samples (as a multiset) whose distance to $S_{\mathrm{priv}}$, $W_{p}(S_{\mathrm{priv}},S_{\mathrm{syn}})$, is minimized. Here $W_{p}$ is the Wasserstein $p$-distance w.r.t. the distance function $d(\cdot,\cdot)$ (see App. B for the definition).
@@
When the distance function $d(\cdot,\cdot)$ is $\ell_{2}$ (in the sample space), DPWA is closely related to DP Clustering (Ghazi et al., 2020; Su et al., 2016; Balcan et al., 2017) and DP Heatmaps (Ghazi et al., 2022). But a direct application of them does not work in our setting; see App. H for more discussions.
@@ 4 Private Evolution (PE)
Inspired by evolutionary algorithms (EA) (Davis, 1987) (App. C), we propose Private Evolution (PE) framework for generating DP synthetic data via APIs. See Fig. 2 for the intuition behind PE. The complete algorithm is in Alg. 1. Below, we discuss the components in detail.
@@
1 $S_{1}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}\right)$
2 for $t\leftarrow 1,\ldots,T$ do
       $histogram_{t}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(S_{\mathrm{priv}},S_{t},\sigma,H\right)$
        // See Alg. 2
       $\mathcal{P}_{t}\leftarrow histogram_{t}/\mathrm{sum}(histogram_{t})$
        // $\mathcal{P}_{t}$ is a distribution on $S_{t}$
       $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}$ samples with replacement from $\mathcal{P}_{t}$
        // $S_{t}^{\prime}$ is a multiset
@@ 3       $S_{t+1}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
1 $histogram\leftarrow[0,\ldots,0]$
2 for $x_{\mathrm{priv}}\in S_{\mathrm{priv}}$ do
3       $i=\arg\min_{j\in\left[n\right]}d{}\left(x_{\mathrm{priv}},z_{j}\right)$
4       $histogram[i]\leftarrow histogram[i]+1$
$histogram\leftarrow histogram+\mathcal{N}\left(0,\sigma I_{n}\right)$
  // Add noise to ensure DP
$histogram\leftarrow\max\left(histogram-H,0\right)$
  // ‘max’, ‘-’ are element-wise
@@ 4       $histogram[i]\leftarrow histogram[i]+1$
In summary, we called the above fitness function DP Nearest Neighbors Histogram. Note that it is not the traditional “histogram” on a continuous space that requires binning. Instead, it is a histogram built on the generated samples: the value of $i$-th bin means the (privatized) number of private samples whose nearest neighbor among the generated ones is the $i$-th sample. See App. A for related work.
@@
Please see App. E for the convergence analysis of PE.
@@ 4.3 Privacy Analysis
Unlike the analysis of DP-SGD which requires complicated DP composition theorems (e.g., Gopi et al. (2021); Mironov (2017)) due to subsampling, PE does not have subsampling steps and therefore the privacy analysis is rather straightforward. The DP guarantee of the unconditional version of PE (Alg. 1) can be reasoned as follows:
@@
For the conditional version of PE (Alg. 3), since it does the unconditional version for each class separately (discussed in § 4.1), adding or removing one sample will only influence the results of one class. For that class, the impact due to the added/removed sample is also bounded, as seen in the privacy analysis above. Therefore, Alg. 3 is also DP. In fact, we can show that the privacy guarantee of Alg. 3 is the same as Alg. 1, and it protects the labels of samples in the same level as DP-SGD (Ghalebikesabi et al., 2023). Please refer to App. D for more details.
@@ 5 Experiments
In § 5.1, we compare PE with SOTA training-based methods on standard benchmarks to understand its promise and limitations. In § 5.2, we present proof-of-concept experiments to show how PE can utilize the power of large foundation models. We did (limited) hyper-parameter tunings in the above experiments; following prior DP synthetic data work (Yu et al., 2021; Ghalebikesabi et al., 2023), we ignore the privacy cost of hyper-parameter tuning. However, as we will see in the ablation studies (§ 5.3 and N), PE stably outperforms SOTA across a wide range of hyper-parameters, and the results can be further improved with better hyper-parameters than what we used. Detailed hyper-parameter settings and more results such as generated samples and their nearest images in the private dataset are in Apps. J, K and L.
@@ 5.1 Comparisons to State-of-the-Art
Public information. We use standard benchmarks (Ghalebikesabi et al., 2023) which treat ImageNet (Deng et al., 2009) as public data. For fair comparisons, we only use ImageNet as public information in PE: (1) Pre-trained model. Unlike the SOTA (Ghalebikesabi et al., 2023) which trains customized diffusion models, we simply use public ImageNet pre-trained diffusion models (pure image models without text prompts) (Nichol & Dhariwal, 2021). RANDOM_API and VARIATION_API are implemented using the same pre-trained model (see App. J). (2) Embedding (Eq. 1). We use ImageNet inception embedding (Szegedy et al., 2016). PE is not sensitive to embedding choice though and we get good results even with CLIP embeddings (Fig. 41 in App. N).
@@
Outline. We test PE on private datasets that are either similar to or differ a lot from ImageNet in § 5.1.1 and 5.1.2. We demonstrate that PE can generate an unlimited number of useful samples in § 5.1.3. We show that PE is computationally cheaper than train-based methods in App. P.
@@ 5.1.2 Large Distribution Shift (ImageNet $\rightarrow$ Camelyon17)
Figure 6: Real and generated images from Camelyon17. More in App. K.
@@
Why PE works under large distribution shifts. Even though the diffusion model is trained on natural images, the support of the generated distribution spans the entire sample space. PE is effective in guiding the model to generate samples from the region that is low-density in the original pre-trained distribution but high-density in the private distribution. See App. K for more explanations.
Limitation. These results demonstrate the effectiveness of PE. But when public models that are similar to private data are not available and when there is enough private data, the traditional training-based methods are still more promising at this point if the privacy-utility trade-off is the only goal. However, given the benefit of API-only assumption and the non-trivial results that PE already got, it is worth further exploiting the potential of PE in future work. Indeed, we expect these results can be improved with further refinement of PE (App. N).
@@ 5.2 More Challenging Benchmarks with Large Foundation Models
Data. Ideally we want to experiment with a dataset that has no overlap with Stable Diffusion’s training data. [^7] The training set of Stable Diffusion is public. However, it is hard to check if a public image or its variants (e.g., cropped, scaled) have been used to produce images in it. Therefore, we resort to our own private data. We take the safest approach: we construct two datasets with photos of the author’s two cats that have never been posted online. Each dataset has 100 512x512 images. Such high-resolution datasets with a small number of samples represent a common need in practice (e.g., in health care), but are challenging for DP synthetic data: to the best of our knowledge, no prior training-based methods have reported results on datasets with a similar resolution or number of samples. The dataset is released at https://github.com/microsoft/DPSDA as a new benchmark. See App. L for all images.
API implementation. We use off-the-shelf Stable Diffusion APIs (see App. L).
Results. We run PE for these two datasets with the same hyperparameters. Fig. 8 show examples of generated images for each of the cat datasets. We can see that Private Evolution correctly captures the key characteristics of these two cats. See App. L for all generated images.
@@
Figure 8: Real & generated images from Cat Cookie (left) and Cat Doudou (right). More in App. L.
@@ 5.3 Ablation Studies
Pre-trained network. Fig. 7 shows the results with two different ImageNet pre-trained networks: one is larger (270M) with ImageNet class labels as input; the other one is smaller (100M) without label input (see App. I for implementation details). In all experiments in § 5.1, we used the 270M network. Two takeaways are: (1) The 270M network (trained on the same dataset) improves the results. This is expected as larger and more powerful models can learn public distributions better. This suggests the potential of PE with future foundation models with growing capabilities. (2) Even with a relatively weak model (100M), PE can still obtain good results that beat the baselines (though with a slower convergence speed), suggesting the effectiveness of PE.
More ablation studies on the lookahead degree $k$, the number of generated samples $N_{\mathrm{syn}}$, the threshold $H$, and the embedding network are in App. N, where we see that PE obtains good results across a wide range of hyper-parameters, and the main results can be improved with better hyper-parameters.
@@ Acknowledgement
This paper is the full version of our previous workshop paper (Lin et al., 2023a).
Appendix A More Related Work
DP data selection. One key component in PE is to use private samples to select similar generated samples (Alg. 2). Prior work studied similar problems in different applications.
In the federated learning setting, Hou et al. (2023) select public datasets that are similar to the clients’ data, and then pre-train the model on the selected datasets before federated fine-tuning. Hong et al. (2022) first cluster public data, and then use clients’ data to select the closest cluster centers (using a similar histogram approach), so that models trained on the selected clusters can have better performance for the clients. In contrast to these works, PE selects generated samples at a sample-level, so that we can improve the generated data in a more fine-grained manner.
Similar to ours, Yu et al. (2023) also conduct sample-level data selection. They select public samples that are similar to the private data for pre-training the model, before DP fine-tuning the model on the private data (Yu et al., 2023). Their data selection rule does not provide guarantees on the distance between the selected samples and the private samples, whereas our selection provides distribution convergence guarantees (App. E).
Nevertheless, given that all these methods deal with DP data selection, they can be interchangeably used in each application. It would be interesting to study such extensions in future work.
Furthermore, we apply such data selection iteratively on the generated data. Together with the use of foundation model APIs, we can generate DP synthetic data, which is not studied in prior work.
Appendix B Definition of Wasserstein Distance
@@
Appendix C A Brief Introduction to Evolutionary Algorithms
@@
Appendix D More Details on Private Evolution
@@
1 $S_{\mathrm{syn}}\leftarrow\emptyset$
2 for $c\in C$ do
3       $private\_samples\leftarrow\left\{x_{i}|(x_{i},y_{i})\in S_{\mathrm{priv}}\text{ and }y_{i}=c\right\}$
4       $S_{1}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}/\left|C\right|\right)$
5       for $t\leftarrow 1,\ldots,T$ do
             $histogram_{t}\leftarrow{\textsf{DP\_NN\_HISTOGRAM}}{}\left(private\_samples,S_{t},\sigma,H\right)$
              // See Alg. 2
             $\mathcal{P}_{t}\leftarrow histogram_{t}/\mathrm{sum}(histogram_{t})$
              // $\mathcal{P}_{t}$ is a distribution on $S_{t}$
             $S_{t}^{\prime}\leftarrow$ draw $N_{\mathrm{syn}}/|C|$ samples with replacement from $\mathcal{P}_{t}$
              // $S_{t}^{\prime}$ is a multiset
@@ 6             $S_{t+1}\leftarrow{\textsf{VARIATION\_API}}{}\left(S_{t}^{\prime}\right)$
Privacy analysis. For ease of understanding the privacy guarantee of Alg. 3, we can consider a modified version of Alg. 3 as in Alg. 4, where we switch the order of the two for loops over $t$ and $c$. Apparently, this modified algorithm gives the same outcome as Alg. 3. The only lines that touch private data are Algs. 4 and 4. The input to these lines is the entire private dataset, and the output is a histogram with size $N_{\mathrm{syn}}$. Same as step 1 in the analysis of Alg. 1 (§ 4.3), each private sample only contributes one vote in the histogram. If we add or remove one sample, the resulting histogram will change by 1 in the $\ell_{2}$ norm. Therefore, the sensitivity of these lines is 1. The following privacy analysis follows exactly the same as steps 2-5 in § 4.3, and therefore, the privacy guarantee of Alg. 3 is the same as Alg. 1. It is important to emphasize that even though Alg. 3 utilizes the labels of the samples directly, the above analysis means that Alg. 3 provides privacy protection to the label assignment of the samples (i.e., the labels each sample has) in the same way as the DP-SGD-fine-tuneing-based algorithms (Ghalebikesabi et al., 2023).
@@
1 $S_{\mathrm{syn}}\leftarrow\emptyset$
2 $S_{1}^{c}\leftarrow{\textsf{RANDOM\_API}}{}\left(N_{\mathrm{syn}}/\left|C\right|\right)$ for each $c\in C$
3 $private\_samples^{c}\leftarrow\left\{x_{i}|(x_{i},y_{i})\in S_{\mathrm{priv}}\text{ and }y_{i}=c\right\}$ for each $c\in C$
4 for $t\leftarrow 1,\ldots,T$ do